OpenStudy (anonymous):

a cell phone company offers a pan for 14.90/month. The first 60 minutes are free and the rate is .10/min after the first 60 minutes. About how many miute can you buy for 30 dollars????

OpenStudy (anonymous):

some one please consider this and help me make an equation

OpenStudy (anonymous):

please to the person looking at this

OpenStudy (anonymous):

Does the 30 dollar include the 14.90 monthly plan cost?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no no yu need to find out the amount ofminutes for 30 bucks and this is algebrc so you need some varabesnd please p in 1 eqatn

OpenStudy (anonymous):

What exactly does it ask?

OpenStudy (anonymous):

amount of minutes of 30 minutes is 60+161=

OpenStudy (anonymous):

I mean make this into a algbraic expression

OpenStudy (anonymous):

with variables and stuff

OpenStudy (anonymous):

make sure you read the question

OpenStudy (anonymous):

x=minutes for a dollar x=10 y= (30-(14.90))x+ 60

OpenStudy (anonymous):

y= (30-14.90))x+ 60

OpenStudy (anonymous):

iranmeah91 u are incoreect the anser is 210 so your equation is wrong

OpenStudy (anonymous):

some one ples give me corect answer

OpenStudy (anonymous):

It is right. You would get 211 x=10 y=(30-14.9)x+60 y=15.1x+60 y=15.1(10)+60 y=151+60 y=211

OpenStudy (anonymous):

umm i said 210 not 211

OpenStudy (anonymous):

Where are you getting your answer from?

OpenStudy (anonymous):

my test that i am doin corections from

OpenStudy (anonymous):

Is your question in full? If not upload a full version exactly as it appear

OpenStudy (anonymous):

Unless you are supposed to round to nearest ten, 211 is coreect

OpenStudy (anonymous):

chill out

OpenStudy (anonymous):

what?